# On the large-time behaviour of affine Volterra processes

### v2

@@ 4. The multidimensional SPDE
The reason is that a component needs to lie on the positive cone if its square root appears somewhere in the SPDE.
We now go through the one-dimensional results of [15, Section 4] and highlight the differences with the multidimensional case presented here. Instead of the spaces $\mathcal{E}^{w}$ introduced in [15, (4.7)], we define for all $n\in\mathbb{N}$,
$\mathcal{E}^{n}:=\left\{\lambda_{0}\in Y^{\star}:\overline{\eta}^{i,n}_{t}\geq 0,\text{ for all }i\in\llbracket 1,\tilde{d}\rrbracket,\;\eta^{i,n}_{t}\in\mathcal{M}(\mathbb{R}_{+},\overline{\mathbb{R}})\text{ for all }i\in\llbracket\tilde{d}+1,d\rrbracket\text{ and }t\geq 0\right\},$
@@
where  $\mathcal{A}^{\star}$ is the linear operator of the form  $\mathcal{A}^{\star}\lambda(\mathrm{d}x)=-x\lambda(\mathrm{d}x)$. Note that $\eta$ satisfies a deterministic equation and the condition for $i\in\llbracket\tilde{d}+1,d\rrbracket$ in the definition of $\mathcal{E}^{n}$ is necessarily satisfied, while we want to restrict the state space of the square-root processes to $\mathbb{R}_{+}$. These definitions allow to define the invariant space $\mathcal{E}:=\cap_{n\in\mathbb{N}}\mathcal{E}^{n}$ and its polar cone :=$\mathcal{E}_{\star}:=\big{\{}y\in Y:\langle y,\lambda\rangle\leq 0\text{ for all }\lambda\in\mathcal{E}\big{\}}$. For $\mathcal{E}$ to be well defined we had to let the coefficients grow with $n$ at a uniform speed across the components, which is a notable difference with the one-dimensional framework of [15] where a single speed $w$ was needed.
[…]
@@
The following is the analogue of [15, Theorem 4.17].
@@
    The SPDE (2.3) admits a unique Markovian solution with values in $\mathcal{E}$ given by a generalised Feller semigroup $(P_{t})_{t\geq 0}$ on $\mathcal{B}^{\varrho}(\mathcal{E})$. The generator $A:\bm{F}\to\mathcal{B}^{\varrho}(\mathcal{E})$ associated to the semigroup $(P_{t})_{t\geq 0}$ reads
    $\mathcal{A}f_{y}(\lambda)=f_{y}(\lambda)\left(\langle\mathcal{A}y,\lambda\rangle+\langle yb(\overline{\lambda}),\nu\rangle+\frac{1}{2}\langle y\sigma^{\top}(\overline{\lambda}),\nu\rangle\langle y\sigma(\overline{\lambda}),\nu\rangle\right)=:f_{y}(\lambda)\mathcal{R}(y,\lambda),$ (4.5)
    where the coefficients are given in (2.5) and $\mathcal{A}$ is the adjoint operator of $\mathcal{A}^{\star}$.  [^1] This is the multidimensional version of the generator given in [15, (4.21)], where there seems to be a mild typo. We have adopted a slightly non-standard notation in that we first define  $\mathcal{A}^{\star}\lambda(\mathrm{d}x)=-x\lambda(\mathrm{d}x)$ and then set $\mathcal{A}$ to be the adjoint operator of $\mathcal{A}^{\star}$ mainly for reasons of being consistent with the existing literature because $\mathcal{A}^{\star}$ is typically considered to act on processes living in $Y^{\star}$.
@@
Before we proceed with the proof of Proposition 4.3, we would like to point out that even though we have calculated the generator $\mathcal{A}f_{y}(\lambda)$ in Proposition 4.3 for Fourier basis elements, we have not calculated the generator for all elements of its domain of definition and very importantly we have not calculated its domain of definition. This is useful to keep in mind as without explicit knowledge of its domain of definition, it can be misleading to use the generator for drawing conclusions about limiting behaviour. As a matter of fact, in the case $K^{i}(t)=\frac{t^{\alpha(i)-1}}{\Gamma(\alpha(i))}\mathrm{e}^{-\delta t}$ with $V_{0}\neq 0$ and $\delta>0$, the calculation in (2.9) shows that the limiting behaviour clearly depends on the initial condition $V_{0}\neq 0$, something that is not reflected in the formula for $\mathcal{A}f_{y}(\lambda)$ for the Fourier basis elements.
@@
The entirety of [15, Section 4] aims at proving [15, Theorem 4.17]. Our result follows from the same lines with a few modifications and remarks that we highlight here. The Lemmas, Assumptions, Remarks, Propositions and Theorems we refer to in this proof are all from [15], unless stated otherwise.
@@
Assumption 4.5 holds. Moreover, Assumption 4.9 is satisfied thanks to our assumption on $K^{i}$ and $R_{i}^{w}$ and Remark 4.10.
@@
The invariant spaces $\mathcal{E}^{n}$ then satisfy all the necessary properties of Proposition 4.8 and Lemma 4.11 because they stay in their convex cone. This yields the generalised Feller property on $\mathcal{B}(\mathcal{E})$ of Theorem 4.13.
@@
We denote the operator $\mathcal{A}^{\star}\lambda(\mathrm{d}x)=-x\lambda(\mathrm{d}x)$ for all $\lambda\in\mathcal{M}(\mathbb{R}_{+},\overline{\mathbb{R}}^{d})$. In the proof of Theorem 4.17, for $i\in\llbracket 1,\tilde{d}\rrbracket$, we replace [15, Equation (4.26)] by
@@
where $N^{ijk,n}_{t}$ is a jump process that jumps by $1/n$ and with intensity $n^{2}\left(\sigma_{ijk}\sqrt{\overline{\lambda}^{k,n}_{t}}+c_{ijk}\right)^{2}$. The same arguments as in the proof of Theorem 4.17 allow to conclude that the limit of $\lambda^{n}$ as $n$ goes to infinity is a generalised Feller process which generator coincides with that of (2.3) that we compute now.
@@
$\lim_{t\downarrow 0}\frac{P_{t}f_{y}(\lambda)-f_{y}(\lambda_{0})}{t}$ $=\lim_{t\downarrow 0}\frac{\lim_{\varepsilon\downarrow 0}\mathbb{E}[F(S^{\varepsilon}_{t})]-f_{y}(\lambda_{0})}{t}$
@@
where the coefficients were given in (2.5). This yields item (i), while item (ii) follows along the same lines as in Theorem 4.17.
@@
$\partial_{t}u(t,\lambda)=\mathcal{A}u(t,\lambda),\qquad u(0,\lambda)=\exp(\langle y_{0},\lambda\rangle).$
@@
$\partial_{t}f_{y_{t}}(\lambda)=f_{y_{t}}(\lambda)\langle\partial_{t}y_{t},\lambda\rangle=f_{y_{t}}(\lambda)\mathcal{R}(y_{t},\lambda)=\mathcal{A}f_{y_{t}}(\lambda),$
@@ 5. Verifying the condition for existence
holds. In this section, we specify the conditions on the parameters of the SPDE that ensure (5.1). Note that under Assumption 2.6 the requirement $\sup_{t\geq 0}\int_{\bm{X}}P_{t}\rho(\lambda)\mathrm{d}\gamma(\lambda)<\infty$ holds automatically because of the deterministic assumption on the initial condition $\lambda_{0}(\mathrm{d}x)=V_{0}\delta_{0}(\mathrm{d}x)$ for some $V_{0}\in D$.
We start with some comments on the possible choices of initial condition $\lambda_{0}$; recall that $\lambda_{0}$ is admissible if it belongs to $\mathcal{E}$. By Assumption 5.2 and Remark 5.3 in [15], $\mathrm{e}^{-r\cdot}\nu\in\mathcal{E}$ for all $r>0$. We can identify different conditions for initial conditions of OU rows or square-root rows. Clearly, for $i\in\llbracket\tilde{d}+1,d\rrbracket$, $\lambda_{0}^{i}$ can be any signed measure on $\overline{\mathbb{R}}$. For $i\in\llbracket 1,\tilde{d}\rrbracket$, by Remark 5.4 in [15] and Example 2.2 in [1], are allowed all $\lambda_{0}\in Y^{\star}$ such that $t\mapsto\int_{0}^{\infty}\mathrm{e}^{-tx}\lambda_{0}(\mathrm{d}x)\in\mathcal{G}_{K}$, where $\mathcal{G}_{K}$ includes in particular Hölder continuous, non-decreasing functions $g$ with $g(0)\geq 0$, and functions of the form $g=V_{0}+K\ast\theta$ such that $\theta(s)\,\mathrm{d}s+V_{0}L(\mathrm{d}s)$ is a non-negative measure, where $V_{0}>0$ and $L$ is the resolvent of first kind of $K$. This yields at least two options:
@@
    $\int_{0}^{\infty}\mathrm{e}^{-xt}\lambda_{0}(\mathrm{d}x)$ $=V_{0}+\Gamma(1-\alpha-\mu)t^{\alpha+\mu-1}=V_{0}+\frac{\Gamma(1-\alpha-\mu)\Gamma(\alpha+\mu)}{\Gamma(\alpha)\Gamma(\mu)}\int_{0}^{t}(t-s)^{\alpha-1}s^{\mu-1}\,\mathrm{d}s,$
@@
[…]
where we used Itô’s isometry. By Lemma 5.2 it is sufficient to prove that the latter is uniformly bounded in $t\geq 0$. Recall that if $k\leq\tilde{d}$ then $\lambda^{k}$ is one-dimensional and, by [3, Lemma 4.2],
@@
<<$\infty$,< ∞ , (5.6)
the last identity following by the integral representation properties of the Mittag-Leffler function, see for example [43, Section 7].
Hence, we indeed get the uniform finiteness in $t\in[0,\infty)$ for $i>\tilde{d}$.
@@
Let us fix $t>0$, then there exist two increasing sequences of sets $(U_{n})_{n\geq 1}$ and $(L_{n})_{n\geq 1}$, both in $\mathfrak{B}(\overline{\mathbb{R}}_{+})$, such that $\lambda_{t}(U_{n})$ is non-negative for all $n\in\mathbb{N}$ and increases towards $\mathfrak{U}(\lambda_{t})$ as $n$ goes to +$\infty$+ ∞, and analogously, $\lambda_{t}(L_{n})$ is non-positive and decreases to $\mathfrak{L}(\lambda_{t})$. We will use the representation

### v3

@@ 4. The multidimensional measure-valued stochastic evolution equation
The reason is that a component needs to lie on the positive cone if its square root appears somewhere in the stochastic evolution equation.
We now go through the one-dimensional results of [17, Section 4] and highlight the differences with the multidimensional case presented here. Instead of the spaces $\mathcal{E}^{w}$ introduced in [17, (4.7)], we define for all $n\in\mathbb{N}$,
[…]
@@
where  $\mathcal{A}^{\star}$ is the linear operator of the form  $\mathcal{A}^{\star}\lambda(\mathrm{d}x)=-x\lambda(\mathrm{d}x)$. Note that $\eta$ satisfies a deterministic equation and the condition for $i\in\llbracket\tilde{d}+1,d\rrbracket$ in the definition of $\mathcal{E}^{n}$ is necessarily satisfied, while we want to restrict the state space of the square-root processes to $\mathbb{R}_{+}$. These definitions allow to define the invariant space $\mathcal{E}:=\cap_{n\in\mathbb{N}}\mathcal{E}^{n}$ and its polar cone $\mathcal{E}_{\star}:=\big{\{}y\in Y:\langle y,\lambda\rangle\leq 0\text{ for all }\lambda\in\mathcal{E}\big{\}}$. For $\mathcal{E}$ to be well defined we had to let the coefficients grow with $n$ at a uniform speed across the components, which is a notable difference with the one-dimensional framework of [17] where a single speed $w$ was needed.
To define the generator of the stochastic evolution equation, we introduce the set $\mathcal{D}:=\{y\in Y:\langle y,\nu\rangle\text{ is well-defined}\}$ and, for each $y\in\mathcal{D}$, the set $\bm{F}$ of Fourier basis elements of the form
@@
The following is the analogue of [17, Theorem 4.17].
@@
    The stochastic evolution equation (2.3) admits a unique Markovian solution with values in $\mathcal{E}$ given by a generalised Feller semigroup $(P_{t})_{t\geq 0}$ on $\mathcal{B}^{\varrho}(\mathcal{E})$. The generator $A:\bm{F}\to\mathcal{B}^{\varrho}(\mathcal{E})$ associated to the semigroup $(P_{t})_{t\geq 0}$ reads
    $Af_{y}(\lambda)=f_{y}(\lambda)\left(\langle\mathcal{A}y,\lambda\rangle+\langle yb(\overline{\lambda}),\nu\rangle+\frac{1}{2}\langle y\sigma^{\top}(\overline{\lambda}),\nu\rangle\langle y\sigma(\overline{\lambda}),\nu\rangle\right)=:f_{y}(\lambda)\mathcal{R}(y,\lambda),$ (4.5)
    where the coefficients are given in (2.5) and $\mathcal{A}$ is the adjoint operator of $\mathcal{A}^{\star}$.  [^1] This is the multidimensional version of the generator given in [17, (4.21)], where there seems to be a mild typo. We have adopted a slightly non-standard notation in that we first define  $\mathcal{A}^{\star}\lambda(\mathrm{d}x)=-x\lambda(\mathrm{d}x)$ and then set $\mathcal{A}$ to be the adjoint operator of $\mathcal{A}^{\star}$ mainly for reasons of being consistent with the existing literature because $\mathcal{A}^{\star}$ is typically considered to act on processes living in $Y^{\star}$.
@@
Before we proceed with the proof of Proposition 4.3, we would like to point out that even though we have calculated the generator $Af_{y}(\lambda)$ in Proposition 4.3 for Fourier basis elements, we have not calculated the generator for all elements of its domain of definition and very importantly we have not calculated its domain of definition. This is useful to keep in mind as without explicit knowledge of its domain of definition, it can be misleading to use the generator for drawing conclusions about limiting behaviour. As a matter of fact, in the case $K^{i}(t)=\frac{t^{\alpha(i)-1}}{\Gamma(\alpha(i))}\mathrm{e}^{-\delta t}$ with $V_{0}\neq 0$ and $\delta>0$, the calculation in (2.9) shows that the limiting behaviour clearly depends on the initial condition $V_{0}\neq 0$.
@@
The entirety of [17, Section 4] aims at proving [17, Theorem 4.17]. Our result follows from the same lines with a few modifications and remarks that we highlight here. The Lemmas, Assumptions, Remarks, Propositions and Theorems we refer to in this proof are all from [17], unless stated otherwise.
@@
[17, Assumption 4.5] holds. Moreover, [17, Assumption 4.9] is satisfied thanks to our assumption on $K^{i}$ and $R_{i}^{w}$ and Remark 4.10.
@@
The invariant spaces $\mathcal{E}^{n}$ then satisfy all the necessary properties of [17, Proposition 4.8 and Lemma 4.11] because they stay in their convex cone. This yields the generalised Feller property on $\mathcal{B}(\mathcal{E})$ of [17, Theorem 4.13].
@@
We recall that $\mathcal{A}^{\star}\lambda(\mathrm{d}x)=-x\lambda(\mathrm{d}x)$ for all $\lambda\in\mathcal{M}(\overline{\mathbb{R}}_{+},\mathbb{R}^{d})$. In the proof of [17, Theorem 4.17], for $i\in\llbracket 1,\tilde{d}\rrbracket$, we replace [17, Equation (4.26)] by
@@
where $N^{ijk,n}_{t}$ is a jump process that jumps by $1/n$ and with intensity $n^{2}\left(\sigma_{ijk}\sqrt{\overline{\lambda}^{k,n}_{t}}+c_{ijk}\right)^{2}$. The same arguments as in the proof of [17, Theorem 4.17] allow to conclude that the limit of $\lambda^{n}$ as $n$ goes to infinity is a generalised Feller process which generator coincides with that of (2.3) that we compute now.
@@
$\lim_{t\downarrow 0}\frac{P_{t}f_{y}(\lambda_{0})-f_{y}(\lambda_{0})}{t}$ $=\lim_{t\downarrow 0}\frac{\lim_{\varepsilon\downarrow 0}\mathbb{E}[F(S^{\varepsilon}_{t})]-f_{y}(\lambda_{0})}{t}$
@@
where the coefficients were given in (2.5). This yields item (i), while item (ii) follows along the same lines as in [17, Theorem 4.17].
@@
$\partial_{t}u(t,\lambda)=Au(t,\lambda),\qquad u(0,\lambda)=\exp(\langle y_{0},\lambda\rangle).$
@@
$\partial_{t}f_{y_{t}}(\lambda)=f_{y_{t}}(\lambda)\langle\partial_{t}y_{t},\lambda\rangle=f_{y_{t}}(\lambda)\mathcal{R}(y_{t},\lambda)=Af_{y_{t}}(\lambda),$
@@ 5. Verifying the condition for existence
holds. In this section, we specify the conditions on the parameters of the stochastic evolution equation that ensure (5.1). We start with some comments on the possible choices of initial condition $\lambda_{0}$; recall that $\lambda_{0}$ is admissible if it belongs to $\mathcal{E}$. By Assumption 5.2 and Remark 5.3 in [17], $\mathrm{e}^{-r\cdot}\nu\in\mathcal{E}$ for all $r>0$. We can identify different conditions for initial conditions of OU rows or square-root rows. Clearly, for $i\in\llbracket\tilde{d}+1,d\rrbracket$, $\lambda_{0}^{i}$ can be any signed measure on $\overline{\mathbb{R}}$. For $i\in\llbracket 1,\tilde{d}\rrbracket$, by Remark 5.4 in [17] and Example 2.2 in [1], are allowed all $\lambda_{0}\in Y^{\star}$ such that $t\mapsto\int_{0}^{\infty}\mathrm{e}^{-tx}\lambda_{0}(\mathrm{d}x)\in\mathcal{G}_{K}$, where $\mathcal{G}_{K}$ includes in particular Hölder continuous, non-decreasing functions $g$ with $g(0)\geq 0$, and functions of the form $g=V_{0}+K\ast\theta$ such that $\theta(s)\,\mathrm{d}s+V_{0}L(\mathrm{d}s)$ is a non-negative measure, where $V_{0}>0$ and $L$ is the resolvent of first kind of $K$. This yields at least two options:
@@
    $\int_{0}^{\infty}\mathrm{e}^{-xt}\lambda_{0}(\mathrm{d}x)$ $=V_{0}+\Gamma(1-\alpha-\mu)t^{\alpha+\mu-1}$
    $=V_{0}+\frac{\Gamma(1-\alpha-\mu)\Gamma(\alpha+\mu)}{\Gamma(\alpha)\Gamma(\mu)}\int_{0}^{t}(t-s)^{\alpha-1}s^{\mu-1}\,\mathrm{d}s,$
@@
$\leq\sum_{k=1}^{\tilde{d}}\mathbb{E}\left[\int_{0}^{\infty}\left(\int_{0}^{t}\mathrm{e}^{-x(t-s)}\beta_{ik}\overline{\lambda}^{k}_{s}\,\mathrm{d}s\right)\nu^{i}(\mathrm{d}x)\right]+\int_{0}^{\infty}\mathbb{E}\left[\int_{0}^{t}\mathrm{e}^{-2x(t-s)}dm\sum_{j=1}^{m}\sum_{k=1}^{\tilde{d}}\sigma_{ijk}^{2}\overline{\lambda}^{k}_{s}\,\mathrm{d}s\right]^{\frac{1}{2}}\nu^{i}(\mathrm{d}x),$
by Itô’s isometry. By Lemma 5.2 it suffices to prove that the latter is uniformly bounded in $t\geq 0$. Recall that if $k\leq\tilde{d}$ then $\lambda^{k}$ is one-dimensional and, by [3, Lemma 4.2],
@@
$<\infty,$ (5.6)
the last identity following by the integral representation properties of the Mittag-Leffler function, see for example [45, Section 7]. Hence, we indeed get the uniform finiteness in $t\in[0,\infty)$ for $i>\tilde{d}$.
@@
Let us fix $t>0$, then there exist two increasing sequences of sets $(U_{n})_{n\geq 1}$ and $(L_{n})_{n\geq 1}$, both in $\mathfrak{B}(\overline{\mathbb{R}}_{+})$, such that $\lambda_{t}(U_{n})$ is non-negative for all $n\in\mathbb{N}$ and increases towards $\mathfrak{U}(\lambda_{t})$ as $n$ goes to $\infty$, and analogously, $\lambda_{t}(L_{n})$ is non-positive and decreases to $\mathfrak{L}(\lambda_{t})$. We will use the representation
@@ 6. Outlook
Acknowledgement
The authors would like to thank the reviewer for a very careful review of the manuscript that identified an issue with the initial version of the result and for allowing us to address it.
Funding
A.P. and A.J. are supported by an EPSRC EP/T032146/1 grant. K.S. is partially supported by the National Science Foundation (DMS 2107856) and Simons Foundation Award 672441.